A ball is projected upwards from a height h above the surface of the earth with velocity v. The time at which the ball strikes the ground is
    (a)    v/g+2hg/√2
    (b)    v/g [1-√(1+2h/g)]
    (c)    v/g [1+√(1+2h/v^2 )]
    (d)    v/g [1+√(v^2+2h/g)]

Grade:upto college level

2 Answers

Radhika Batra
10 years ago

(c)

 

        Using vector form of kinematic equation

 

        –h = uT –  gT2

 

        T =

Radhika Batra
10 years ago

(c)

    Using vector form of kinematic equation

    –h = uT – 1/2 gT2

    T = (v+√(v^2+2gh))/g=v/g [1+√(1+2gh/v^2 )]